OpenStudy (anonymous):

What are the exact solutions of x^(2)+3x+1=0?

OpenStudy (anonymous):

Use the quadratic formula: \[ax^2+bx+c=0~\Rightarrow~x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (uri):

^ this and a:1 b:3 c:-1 If you haev to solve it.